# Denoising Diffusion Probabilistic Models for Coastal Inundation Forecasting

### v1

@@ Appendix C Inundation Scenario Queries
Once Diff-Flood is trained, we can use it to answer inundation scenario-based queries relevant to policymakers. In this Section, we discuss a few of such queries.
@@
Input: Query area $A$ as a Polygon, Flooding level threshold $d$, Prediction Length $T$, Trained Diff-Flood Model.
1 $C_{A}\leftarrow$ Set of grid cells that overlap with $A$.
@@ 6       $probability(P_{\leq d})\leftarrow num\_scenarios(P_{\leq d})/M$
    Since we require the probability of flooding level being above $d$ units in area $A$, our desired probability value is $(1-p)$.
@@
It is important to note that, we can have more complex queries. For instance, a more refined version of Query 1 can be: what is the probability that at-least half of the region $A$ will have flooding level above $d$ units in the next $T$ hours? The key idea for answering such queries is that, we have a sampling method in the form a trained Diff-Flood model and we can use it to sample inundation scenarios from the distribution of possible scenarios. We use the sampled scenarios to calculate our desired probability value.
Appendix D Hyper-parameters
@@
Learning rate update step size 5 epochs
@@
NO. of scenarios generated for validation 2
NO. of scenarios generated for test 8
@@
Appendix E Patch Visualization

### v2

@@ Appendix E Inundation Scenario Queries
Once Diff-Sparse is trained, we can use it to answer inundation scenario-based queries relevant to policymakers. In this Section, we discuss a few of such queries.
@@
Input: Query area AA as a Polygon, Flooding level threshold $d$, Prediction Length $T$, Trained Diff-Sparse Model.
1 $C_{A}\leftarrow$ Set of grid cells that overlap with AA.
@@ 6 $probability(P_{\leq d})\leftarrow num\_scenarios(P_{\leq d})/M$
    Since we require the probability of flooding level being above $d$ units in area $A$, our desired probability value is (1p)(1-p).
@@
It is important to note that, we can have more complex queries. For instance, a more refined version of Query 1 can be: what is the probability that at-least half of the region $A$ will have flooding level above $d$ units in the next $T$ hours? The key idea for answering such queries is that, we have a sampling method in the form a trained Diff-Sparse model and we can use it to sample inundation scenarios from the distribution of possible scenarios. We use the sampled scenarios to calculate our desired probability value.
Appendix F Reproducibility Checklist Additional Information
Our source code for Diff-Sparse and comparison with baseline methods are provided with the supplementary materials. The TideWatch inundation dataset will be provided upon request.
Appendix G Hyper-parameters
@@
Learning rate update policy Reduce on Plateau.
Learning rate update patience 3 epochs.
@@
# of scenarios generated for validation 2
# of scenarios generated for test 8
@@
Appendix H Patch Visualization